List all the factors of x2 + 3x + 2
(x + 1)
(x + 2)
(x + 3)
(x - 1)
(x - 2)
(x - 3)

Answer by checkley71(8403)

x^2+3x+2=0
(x+2)(x+1)=0
(x+2) solution
(x+1) solution